A CBS News poll conducted June 10 and 11, 2006, among a nationwide random sample of 651 adults, asked those adults about their party affiliation (Democrat, Republican or none) and their opinion of how the US economy was changing ("getting better," "getting worse" or "about the same"). The results are shown in the table below.

better same worse
Republican 38 104 44
Democrat 12 87 137
none 21 90 118



Express your answers as a decimal and round to the nearest 0.001 (in other words, type 0.123, not 12.3% or 0.123456).

If we randomly select one of the adults who participated in this study, compute:

P(affiliated with neither party) =

P(worse) =

P(worse|affiliated with neither party) =

P(affiliated with neither party|worse) =

P(affiliated with neither party and worse) =
Answer #1
better same worse total
Republican 38 104 44 186
Democrat 12 87 137 236
none 21 90 118 229
total 71 281 299 651

1)

P(affiliated with neither party) = 229/651 =0.352

2)

P(worse) = 299/651 =0.459

3)

P(worse|affiliated with neither party) = 118/229 =0.515

4)

P(affiliated with neither party|worse) = 118/299 =0.395

5)

P(affiliated with neither party and worse) = 118/651 =0.181